Status Report

NASA ISS On-Orbit Status 2 August 2008

By SpaceRef Editor
August 3, 2008
Filed under , , ,
NASA ISS On-Orbit Status 2 August 2008

All ISS systems continue to function nominally, except those noted previously or below. Saturday – mostly off duty for CDR Volkov, FE-1 Kononenko & FE-2 Chamitoff.

Gregory Chamitoff had Part 1 of his third session with the NASA/JSC experiment NUTRITION w/Repository, for blood collection only, for which he had to forego exercising and food intake since yesterday for eight hours. Later today, the FE-2 will also set up the equipment for the 24-hour urine collections which start with the first void early tomorrow morning. [After performing self-phlebotomy, i.e., drawing blood samples (from an arm vein), the samples were first allowed to coagulate in the Repository for 20-30 minutes, then spun in the HRF RC (Human Research Facility/Refrigerated Centrifuge) and finally placed in MELFI (Minus-Eighty Laboratory Freezer for ISS). No thruster activity was allowed during the blood drawing. The RC was later powered off after a temperature reset to limit wear on the compressor, and cleaned.]

The crew performed the regular weekly three-hour task of thorough station cleaning. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the house cleaning, the crew also conducted regular maintenance inspection & cleaning on fan grilles in the FGB (TsV2), SM (VPkhO, VPrK, FS5, FS6 & FS9), DC1 (V3) screens and PF1/PF2 dust filters, as well as in the COL (Columbus Orbital Laboratory).

The FE-1 also checked up on the Russian POTOK-150MK (150 micron) air filter unit of the SM’s SOGS air revitalization subsystem, gathering weekly data on Total Operating Time & “On” durations for reporting to TsUP-Moscow. Later, CDR Volkov temporarily powered down the POTOK air filtration system for the periodic cleaning of its pre-filter, using the vacuum cleaner with narrow-slit nozzle attachment.

Gregory filled out the regular FFQ (Food Frequency Questionnaire), his eighth, on the MEC (Medical Equipment Computer). [On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.]

Chamitoff also had about half an hour reserved to work on the CMRS (Crew Medical Restraint System), stowed in the CHeCS (Crew Health Care Systems) rack, performing the periodic checkout and inspection of the system for upcoming standard CMO (Crew Medical Officer) proficiency training. [The FE-2 inspected the CMRS for cracks in the board and/or metal fastener exposed on top of CMRS (found on the ground units), either of which could provide a high-voltage defibrillation ground path from the patient to ISS structure. The board-like CMRS allows strapping down a patient on the board with a harness for medical attention by the CMO who is also provided with restraints around the device. The device can be secured to the ISS structure within two minutes to provide a patient restraint surface for performing emergency medical procedures, such as during ACLS (Advanced Cardiac Life Support). It can also be used to transport a patient between the station and the Orbiter middeck. It isolates the crew and equipment electrically during defibrillations and pacing electrical discharges, accommodates the patient in the supine zero-G positions, provides cervical spine stabilization and, for a three-person crew, can also restrain two CMOs during their delivery of medical care.]

At ~10:00am EDT, the crewmembers conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

For today’s VolSci (Voluntary Weekend Science) program featuring SHERE (Shear History Extensional Rheology Experiment), FE-2 Chamitoff –

  • Activated the MSG (Microgravity Science Glovebox),
  • Powered on the SHERE hardware,
  • Accessed the CGBA (Commercial Generic Bioprocessing Apparatus) to install the SHERE FM (Fluid Module) #25,
  • Supported the first SHERE experiment run (Test Point 3),
  • Transferred the module with the fluid sample,
  • Installed FM #24 for the second experiment run (Test Point 13), ,
  • Removed the FM from the CGBA, followed by SHERE data transfer,
  • Turned off the SHERE/CGBA equipment,
  • Transferred the data files to the MSG laptop, and
  • Powered down the MSG.

.Later, Chamitoff completed a run with the MedOps experiment WinSCAT (Windows Spaceflight Cognitive Assessment Tool), his second onboard session, by logging in on the MEC (Medical Equipment Computer) and performing the psychological evaluation exercise on the laptop-based WinSCAT experiment. [WinSCAT is a monthly time-constrained questionnaire test of cognitive abilities, routinely performed by astronauts aboard the ISS every 30 days before or after the PHS (periodic health status) test or on special CDR’s, crewmembers or flight surgeons request.]

Kononenko performed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the Service Module (SM). [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

The three crewmembers conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), RED resistive exercise device (CDR, FE-1, FE-2), and VELO bike with bungee cord load trainer (CDR, FE-1). Later, Oleg transferred the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure
on the HRM storage medium (done six times a week).

At ~6:35am EDT, the FE-1 had his weekly PFC (Private Family Conference) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop).

Weekly Science Update (Expedition Seventeen — Week 15)

3-D SPACE: Third session by Greg Chamitoff has been successfully performed on 7/30.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in the FGB module.

ANITA (Analyzing Interferometer for Ambient Air): Continuing.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Samples returned on 1J.

CSI-2/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): “On 7/31, we had a successful activation of Silicate Garden Hab Side 2 in the afternoon and the crystals looked great at the start of the growth period. Images are being taken once per 90 seconds and will be decreased to once per 5 minutes at 7:00pm EDT today.”

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): On 7/16, EuTEF encountered a MIL-Bus error (this happened similarly in the past). The platform has to be power-cycled and the entire EuTEF facility re-initiated. The instruments have been re-activated with some delay due to some minor mismatch in the sequence of commands sent from ground. Although this recurrent DHPU (Data Handling Processing Unit) problem does not prevent EuTEF to function, ground teams are working hard to characterize it and find a fix. A DHPU software patch is currently being developed to fix the link error issues encountered with DEBIE-2 and FIPEX instruments.– DEBIE-2: Instrument could not be restarted immediately after the DHPU problem on 7/27, and was finally re-activated on 7/28 for 2 x 24hrs science runs. Generic status: link error still in work. However, a work-around allows for regular science data acquisition using an on-board IOP (Instrument Operations Procedure), but not yet in conjunction with any FIPEX IOP – a final software patch, which will enhance the present instrument capabilities, is under finalization on ground;– DOSTEL: On-going science acquisition;– EuTEMP: Currently inactive as planned;– EVC: Inactive this week;– EXPOSE: On-going science acquisition;– FIPEX: Instrument re-activated after DHPU problem on 7/27. Science acquisition since 7/30; — MEDET: Instrument re-activated after DHPU problem on 7/27. Science acquisition since 7/30; – PLEGPAY: Currently in READY mode, no science data acquisition on-going;–TRIBOLAB: The instrument is currently in Stand-By Mode. The Ball-Bearing experiment would be restarted towards mid-Aug. Analysis is on-going on ground.

FSL (Fluid Science Laboratory): FSL is back but off. Further FSL commissioning activities remain to be done (see following item).

GEOFLOW: Thanks, Greg. for inserting the GEOFLOW Experiment Container on 7/28. However, it was not possible to activate FSL, as some laser switch inhibit did not seem to be positively engaged. That prevents to power on GEOFLOW and later on to proceed with check-out routines and finally first science runs. Ground teams are finalizing a corrective procedure for the crew.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Planned.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer, ESA): In progress.


PADLES (Passive Dosimeter for Lifescience Experiment in Space): In progress.

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SHERE (Shear History Extensional Rheology Experiment): The SHERE team thanks Greg for selecting SHERE for Voluntary Science checkout activities performed on GMT 202. The checkout tests were a great success and showed that the SHERE hardware is functioning properly. The SHERE team looks forward to working with Greg again tomorrow, GMT 208, to perform the SHERE Dry Run. The recent fit problem with the Fluid Module Stowage Tray is being worked and will not affect the Dry Run.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): The HRP team thanks Greg for completing his first week of Sleep logs and his Actiwatch download initialization session. HRP is currently targeting the week of 8/4 for Greg’s next week of Sleep logging and the week of 8/11 for his next Actiwatch download. Additional Sleep logging on non-targeted weeks is above and beyond and greatly appreciated by the PI.

SOLAR (Solar Monitoring Observatory): On 7/28, SOLAR experienced a spontaneous reboot, which was recovered on 7/29. It looks like SOLAR does not communicate anymore with the Columbus LAN network when under Columbus LAN switch#2 (CLSW#2) configuration. With the platform connected to the Columbus LAN switch #1 (CLSW#1), the SOLAR team is confident that everything works now for the next Sun observation window which starts on 7/31.– SOVIM: Instrument confirmed fully functional; awaiting the Sun; — SOLSPEC: Instrument confirmed fully functional; awaiting the Sun;– SOLACES: Instrument confirmed fully functional; awaiting the Sun.

SOLO: Planned.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The four ECs (Experiment Containers) of Rotor B are yet to be retrieved by the crew from the blocked Rotor B. BIOLAB could not be activated from ground yet; pending resolution of the Smoke Detector issue. A troubleshooting plan has been developed to recover all the BIOLAB functionalities. The first step of the BIOLAB recovery plan was to exchange an ESEM Power Board and this activity has been successfully performed on 7/25. On 7/26, BIOLAB was activated from ground for about 1 hour, in order to get telemetry about the Incubator Smoke Detector sensor. This test was positive, but engineering team would like to gather additional data during a longer activation test to confirm full functionality of the sensor. Next steps will follow, starting with the BIOLAB Rotor A bellow test (ground only) and later on the remaining WAICO-#1 ECs will be disposed and replaced by six Reference ECs on Rotor B. Planning for this last activity is currently TBD.

CEO (Crew Earth Observations): Through 7/15 the ground has received a total of 3,734 frames of CEO images for review and cataloging. “We have received some additional imagery this past week, but none with camera times corresponding to our CEO target request times. Most of these appear to be in South America and of good quality. One of your more striking Polar Mesopheric Cloud (PMC) photos, acquired last week, in a pass over central Asia will be published on NASA/GSFC’s Earth Observatory webpage. Your PMC photos a welcome addition to our growing collection of images relevant to ISS program support of the International Polar Year (IPY). Thanks for your diligence in acquiring them!”

CEO (Crew Earth Observations) photo targets uplinked for today were Red River Basin, TX (documenting the land use change and water levels of this meandering river boundary between Texas and Oklahoma, using the long lens settings and map along the river westward from Lake Texoma. This was a fair-weather nadir pass in mid-afternoon with an approach from the SW), Santa Barbara Coast, California (this target is a Long Term Ecological Research [LTER] site centered near Santa Barbara, CA. On this nadir, late-afternoon pass, clear skies were expected over the land areas as ISS approached from the SW. Using the long lens settings for a detailed mapping of the Santa Inez Mountains from Lompoc in the west to just north of Ventura to the east), and Slate Islands Impact Crater (this 450 million year old impact is about 30-km in diameter and is revealed as a small cluster of island near the north coast of Lake Superior. As ISS tracked ENE- ward to the north of Lake Superior in late afternoon, Greg was to look for the target just right of track, using the long lens settings for detail).

CEO photography can be studied at this “Gateway” website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

Significant Events Ahead (all dates Eastern Time, some changes possible.):
08/13/08 — ATV Reboost
08/30/08 — Progress M-64/29P undocking, from FGB nadir
09/05/08 — ATV1 undocking, from SM aft port (loiter until ~9/25 for nighttime reentry/observation)
09/10/08 — Progress M-65/30P launch
09/12/08 — Progress M-65/30P docking (SM aft port)
10/01/08 — NASA 50 Years (official)
10/08/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
10/11/08 — Progress M-65/30P undocking (from SM aft port)
10/12/08 — Soyuz TMA-13/17S launch
10/14/08 — Soyuz TMA-13/17S docking (FGB nadir port)
10/23/08 — Soyuz TMA-12/16S undocking (DC1 nadir)
11/10/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC
11/12/08 — STS-126/Endeavour/ULF2 docking
11/20/08 — ISS 10 Years
11/26/08 — Progress M-66/31P launch
11/28/08 — Progress M-66/31P docking
02/10/09 — Progress M-67/32P launch
02/12/09 — Progress M-67/32P docking
02/12/09 — STS-119/Discovery/15A launch – S6 truss segment
03/25/09 — Soyuz TMA-14/18S launch
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
07/30/09 — STS-128/Atlantis/17A – MPLM(P), last crew rotation
05/27/09 — Six-person crew on ISS (following Soyuz 19S docking, May ’09)
10/15/09 — STS-129/Discovery/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A – MPLM(P)
04/08/10 — STS-132/Discovery/ ULF4 – ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4 (contingency).

SpaceRef staff editor.